# Linear Mode Connectivity in Differentiable Tree Ensembles

### v1

@@ 3.2 Matching Strategy
Here, we propose a matching strategy for binary trees. When considering invariances, it is necessary to compare multiple functionally equivalent trees and select the most suitable one for achieving LMC. Although comparing tree parameters is a straightforward approach, since the contribution of all the parameters in a tree is not equal, we apply weighting for each node for better matching. By interpreting a tree as a rule set with shared parameters as shown in Figure 3, we determine the weight of each splitting node by counting the number of leaves to which the node affects. For example, in the case of the left example in Figure 3, the root node affects eight leaves, nodes at depth $2$ affect four leaves, and nodes at depth $3$ affect two leaves. This strategy can apply to even trees other than perfect binary trees. For example, in the right example of Figure 3, the root node affects four leaves, a node at depth $2$ affects three leaves, and a node at depth $3$ affects two leaves.
In this paper, we employ the LAP, which is used as a standard benchmark [4] for matching algorithms. The procedures for AM and WM are as follows. Detailed algorithms (Algorithms 1 and 2) are described in Section A in the supplementary material.
  • Activation Matching (Algorithm 1). In trees, there is nothing that directly corresponds to the activations in neural networks. However, by treating the output of each individual tree as an activation value of a neural network, it is possible to optimize the permutation of trees while examining their output similarities. Regarding subtree flip and splitting order invariances, it is possible to find the optimal pattern from all the possible patterns of flips and changes in the splitting order. Since the tree-wise output remains unchanged, the similarity between each tree, generated by considering additional invariances, and the target tree is evaluated based on the inner product of parameters while applying node-wise weighting.
  • Weight Matching (Algorithm 2). Similar to AM, WM also involves applying weighting while extracting the optimal pattern by exploring possible flipping and ordering patterns. Although it is necessary to solve the LAP multiple times for each layer in MLPs [4], tree ensembles require only a single run of the LAP since there are no layers.
The time complexity of solving the LAP is $\mathcal{O}(M^{3})$ using a modified Jonker-Volgenant algorithm without initialization [26], implemented in SciPy [27], where $M$ is the number of trees. If only considering tree permutation, this process needs to be performed only once in both WM and AM. However, when considering additional invariances, we need to solve the LAP for each pattern generated by considering these additional invariances. In a non-oblivious perfect binary tree with depth $D$, there are $12^{D}-$ 1 splitting nodes, leading to $2^{2^{D}-1}$ possible combinations of sign flips. Additionally, in the case of oblivious trees, there are $D!$ different patterns of splitting order invariance. Therefore, for large values of $D$, conducting a brute-force search becomes impractical.
3.3 Architecture-dependency of the Invariances
In previous subsections, tree architectures are fixed to perfect binary trees as they are most commonly and practically used in soft trees. However, tree architectures can be flexible as we have shown in the right example in Figure 3, and here we show that we can specifically design tree architecture that has neither the subtree flip nor splitting order invariances. This allows efficient matching as considering such two invariances is computationally expensive.
Table 1: Invariances inherent to each model architecture.
Perm Flip Order
Non-Oblivious Tree $\checkmark$ $\checkmark$ $\times$
Oblivious Tree $\checkmark$ $\checkmark$ $\checkmark$
Decision List $\checkmark$ ()($\checkmark$)( ✓ ) $\times$
Decision List (Modified) $\checkmark$ $\times$ $\times$
4 Experiment
4.1 Setup
Datasets.  In our experiments, we employed Tabular-Benchmark [28], a collection of tabular datasets suitable for evaluating tree ensembles. Details of datasets are provided in Section B in the supplementary material. As proposed in [28], we randomly sampled $10,000$ instances for train and test data from each dataset. If the dataset contains fewer than $20,000$ instances, they are randomly divided into halves for train and test data. We applied quantile transformation to each feature and standardized it to follow a normal distribution.
Resources.  All experiments were conducted on a system equipped with an Intel Xeon E5-2698 CPU at 2.20 GHz, 252 GB of memory, and Tesla V100-DGXS-32GB GPU, running Ubuntu Linux (version 4.15.0-117-generic).
4.2 Results for Perfect Binary Trees
Figure 5 shows how the barrier between two perfect binary tree model pairs changes in each operation. The vertical axis of each plot in Figure 5 shows the averaged barrier over datasets for each considered invariance. The results for both the oblivious and non-oblivious trees are plotted separately in a vertical layout. The panels on the left display the results when the depth $D$ of the tree varies, keeping $M=256$ constant. The panels on the right show the results when the number of trees $M$ varies, with $D$ fixed at $2$. For both oblivious and non-oblivious trees, we observed that the barrier significantly decreases as the considered invariances increase. Focusing on the test data results, after accounting for various invariances, the barrier is nearly zero, indicating that LMC has been achieved.
Figure 6: Interpolation curves of test accuracy for oblivious trees on 16 datasets from Tabular-Benchmark [28]. Two model pairs are trained with on the same dataset. The error bars show the standard deviations of $5$ executions. We used $M=256$ trees with a depth $D=2$.
In particular, the difference between the case of only permutation and the case where additional invariances are considered tends to be larger in the case of AM. This is because parameter values are not used during the rearrangement of the tree in AM. Additionally, it has been observed that the barrier increases as trees become deeper, and the barrier decreases as the number of trees increases. These behaviors correspond to the changes observed in neural networks when the depth varies or when the width of hidden layers increases [3, 4]. Figure 6 shows interpolation curves when using AM in oblivious trees with $D=2$ and $M=256$. Other detailed results, such as performance for each dataset, are provided in Section C in the supplementary material.
Figure 7: Interpolation curves of test accuracy for oblivious trees on 16 datasets from Tabular-Benchmark [28]. Two model pairs are trained on split datasets with different class ratios. The error bars show the standard deviations of $5$ executions. We used $M=256$ trees with a depth $D=2$.
Furthermore, we conducted experiments with split data following the protocol in [4, 30], where the initial split consists of randomly sampled 80% negative and 20% positive instances, and the second split inverts these ratios. There is no overlap between the two split datasets. We trained two model pairs using these separately split datasets and observed an improvement in performance by interpolating their parameters. Figure 7 illustrates the interpolation curves under AM in oblivious trees with parameters $D=2$ and $M=256$. We can observe that considering additional invariances improves performance after interpolation. Note that the data split is configured to remain consistent even when the training random seeds differ. Detailed results for each dataset using WM or AM are provided in Section C of the supplementary material.
MLP
Barrier
Depth Naive Perm [4] Accuracy Size
1 8.755 ± 0.877 0.491 ± 0.062 76.286 ± 0.094 12034
2 15.341± 1.125 2.997 ± 0.709 75.981 ± 0.139 77826
3 15.915 ± 2.479 5.940 ± 2.153 75.935 ± 0.117 143618
Non-Oblivious Tree
Barrier
Depth Naive Perm Ours Accuracy Size
1 8.965 ± 0.963 0.449 ± 0.235 0.181 ± 0.078 76.464 ± 0.167 12544
2 6.801 ± 0.464 0.811 ± 0.333 0.455 ± 0.105 76.631 ± 0.052 36608
3 5.602 ± 0.926 1.635 ± 0.334 0.740 ± 0.158 76.339 ± 0.115 84736
Oblivious Tree
Barrier
Depth Naive Perm Ours Accuracy Size
1 8.965 ± 0.963 0.449 ± 0.235 0.181 ± 0.078 76.464 ± 0.167 12544
2 7.881 ± 0.866 0.918 ± 0.092 0.348 ± 0.172 76.623 ± 0.042 25088
3 7.096 ± 0.856 1.283 ± 0.139 0.484 ± 0.049 76.535 ± 0.063 38656
Table 2 compares the average test barriers of an MLP with a ReLU activation function, whose width is equal to the number of trees, $M=256$. The procedure for MLPs follows that described in Section 4.1. The permutation for MLPs is optimized using the method described in [4]. Since [4] indicated that WM outperforms AM in neural networks, WM was used for the comparison. Overall, tree models exhibit smaller barriers compared to MLPs while keeping similar accuracy levels. It is important to note that MLPs with $D>1$ tend to have more parameters at the same depth compared to trees, leading to more complex optimization landscapes. Nevertheless, the barrier for the non-oblivious tree at $D=3$ is smaller than that for the MLP at $D=2$, even with more parameters. Furthermore, at the same depth of $D=1$, tree models have a smaller barrier. Here, the model size is evaluated using $F=44$, the average input feature size of 16 datasets used in the experiments.
In Section 3.2, we have shown that considering additional invariances for deep perfect binary trees is computationally challenging, which may suggest developing heuristic algorithms for deep trees. However, we consider it is rather a low priority, supported by our observations that the barrier tends to increase as trees deepen even if we consider invariances. This trend indicates that deep models are fundamentally less important for model merging considerations. Furthermore, deep perfect binary trees are rarely used in practical scenarios. [12] have demonstrated that generalization performance degrades with increasing depth in perfect binary trees due to the degeneracy of the Neural Tangent Kernel (NTK) [31]. This evidence further supports the preference for shallow perfect binary trees, and increasing the number of trees can enhance the expressive power while reducing barriers.
4.3 Results for Decision Lists
Table 3: Barriers averaged for 16 datasets under WM with $D=2$ and $M=256$.
Train Test
Barrier Barrier
Architecture Naive Perm Ours Accuracy Naive Perm Ours Accuracy
Table 4: Barriers averaged for 16 datasets under AM with $D=2$ and $M=256$.
Train Test
Barrier Barrier
Architecture Naive Perm Ours Accuracy Naive Perm Ours Accuracy
Figure 8 illustrates barriers as a function of depth, considering only permutation invariance, with $M$ fixed at $256$. In this experiment, we have excluded non-oblivious trees from comparison as the number of their parameters exponentially increases as trees deepen, making them infeasible computation. Our proposed modified decision lists reduce the barrier more effectively than both oblivious trees and the original decision lists. However, barriers of the modified decision lists are still larger than those obtained by considering additional invariances with perfect binary trees. Tables LABEL:tbl:wm_models and LABEL:tbl:am_models show the averaged barriers for 16 datasets, with $D=2$ and $M=256$. Although barriers of modified decision lists are small when considering only permutations (Perm), perfect binary trees such as oblivious trees with additional invariances (Ours) exhibit smaller barriers, which supports the validity of using oblivious trees as in [9, 11]. To summarize, when considering the practical use of model merging, if the goal is to prioritize efficient computation, we recommend using our proposed decision list. Conversely, if the goal is to prioritize barriers, it would be preferable to use perfect binary trees, which have a greater number of invariant operations that maintain the functional behavior.
5 Conclusion
We have presented the first investigation of LMC for soft tree ensembles. We have identified additional invariances inherent in tree architectures and empirically demonstrated the importance of considering these factors. Achieving LMC is crucial not only for understanding the behavior of non-convex optimization from a learning theory perspective but also for implementing practical techniques such as model merging. By arithmetically combining parameters of differently trained models, a wide range of applications such as task-arithmetic [32], including unlearning [33] and continual-learning [34], have been explored. Our research extends these techniques to soft tree ensembles that began training from entirely different initial conditions. We will leave these empirical investigations for future work.
This study provides a fundamental analysis of ensemble learning, and we believe that our discussion will not have any negative societal impacts.
Appendix A Detailed Algorithms
We present pseudo-code of algorithms for activation matching (Algorithm 1) and weight matching (Algorithm 2). In these algorithms, if there is only one possible pattern for $U\in\mathbb{N}$, which represents the number of possible operations, and the corresponding operation does nothing in particular, it becomes equivalent to simply considering tree permutations.
1 ActivationMatching($\boldsymbol{\Theta}_{A}\in\mathbb{R}^{M\times P_{\text{{Tree}}}}$, $\boldsymbol{\Theta}_{B}\in\mathbb{R}^{M\times P_{\text{{Tree}}}}$, $\boldsymbol{x}_{\text{{sampled}}}\in\mathbb{R}^{F\times N_{\text{{sampled}}}}$)
@@ 18      Initialize operation indices $\boldsymbol{q}\in\mathbb{N}^{M}$
Algorithm 1 Activation matching for soft trees
1 WeightMatching($\boldsymbol{\Theta}_{A}\in\mathbb{R}^{M\times P_{\text{{Tree}}}}$, $\boldsymbol{\Theta}_{B}\in\mathbb{R}^{M\times P_{\text{{Tree}}}}$)
@@ 3       Initialize similarity matrix for each operation $\boldsymbol{S}\in\mathbb{R}^{U\times M\times M}$
                    // $\boldsymbol{\theta}\in\mathbb{R}^{P_{\text{{Tree}}}}$: Adjusted tree-wise parameters
@@
Algorithm 2 Weight matching for soft trees
@@
Here, we describe the specifications of the notations and functions used in Algorithms 1 and 2. In Section 2.1, $\boldsymbol{\Theta}_{A}$ and $\boldsymbol{\Theta}_{B}$ are initially defined as vectors. However, for ease of use, in Algorithms 1 and 2, $\boldsymbol{\Theta}_{A}$ and $\boldsymbol{\Theta}_{B}$ are represented as matrices of size $\mathbb{R}^{M\times P_{\text{{Tree}}}}$, where $P_{\text{{Tree}}}$ denotes the number of parameters in a single tree. Multidimensional array elements are accessed using square brackets [$\cdot$]. For example, for $\boldsymbol{G}\in\mathbb{R}^{I\times J}$, $\boldsymbol{G}[i]$ refers to the $i$th slice along the first dimension, and $\boldsymbol{G}[:,j]$ refers to the $j$th slice along the second dimension, with sizes $\mathbb{R}^{J}$ and $\mathbb{R}^{I}$, respectively. Furthermore, it can also accept a vector $\boldsymbol{v}\in\mathbb{N}^{l}$ as an input. In this case, $\boldsymbol{G}[\boldsymbol{v}]\in\mathbb{R}^{l\times J}$. The Flatten function converts multidimensional input into a one-dimensional vector format. As the LinearSumAssignment function, $\operatorname{scipy.optimize.linear\_sum\_assignment}$ [^3] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.linear_sum_assignment.html is used to solve the LAP. In the AdjustTree function, the parameters of a tree are modified according to the $u$th pattern among the enumerated $U$ patterns. Additionally, in the Weighting function, parameters are multiplied by the square root of their weights defined in Section 3.2 to simulate the process of assessing a rule set. If the first argument for the UpdateBestOperation function, the input inner product, is larger than any previously input inner product values, then $u^{\prime}$ is updated with $u$, the second argument. If not, $u^{\prime}$ remains unchanged.
Appendix B Dataset

### v2

@@ 3.2 Matching Strategy
When considering subtree flip invariance and splitting order invariance, it is necessary to compare multiple functionally equivalent trees and select the most suitable one for achieving LMC. Although comparing tree parameters is a straightforward approach, since the contribution of all the parameters in a tree is not equal, we apply appropriate weighting for each node. By interpreting a tree as a rule set with shared parameters as shown in Figure 4, we determine the weight of each splitting node by counting the number of leaves to which the node affects. For example, in the case of the example in the left-hand side of Figure 4, the root node affects eight leaves, nodes at depth $2$ affect four leaves, and nodes at depth $3$ affect two leaves. This strategy can apply to even trees other than perfect binary trees. For example, in the right example of Figure 4, the root node affects four leaves, a node at depth $2$ affects three leaves, and a node at depth $3$ affects two leaves.
1 ActivationMatching($\boldsymbol{\Theta}_{A}\in\mathbb{R}^{M\times P}$, $\boldsymbol{\Theta}_{B}\in\mathbb{R}^{M\times P}$, $\boldsymbol{x}_{\text{{sampled}}}\in\mathbb{R}^{F\times N_{\text{{sampled}}}}$)
@@ 18      Initialize operation indices $\boldsymbol{q}\in\mathbb{N}^{M}$
Algorithm 1 Activation matching for soft tree ensembles
1 WeightMatching($\boldsymbol{\Theta}_{A}\in\mathbb{R}^{M\times P}$, $\boldsymbol{\Theta}_{B}\in\mathbb{R}^{M\times P}$)
@@ 3       Initialize similarity matrix for each operation $\boldsymbol{S}\in\mathbb{R}^{U\times M\times M}$
                    // $\boldsymbol{\theta}\in\mathbb{R}^{P}$: Adjusted tree-wise parameters
@@
Algorithm 2 Weight matching for soft tree ensembles
Using the weighting operation described above, we present the straightforward matching procedure in Algorithms 1 and 2. We performed an exhaustive search to explore all patterns with subtree flip invariance and splitting order invariance, while handling tree permutation invariance with the LAP. We treat the output of each individual tree like the activation value of a neural network in the case of AM. Note that although it is necessary to solve the LAP multiple times for each layer in MLPs to perform coordinate descent [4], tree ensembles require only a single run of the LAP since there is no concept of intermediate layers.
@@
Notations used in Algorithms 1 and 2.  Multidimensional array elements are accessed using square brackets [$\cdot$]. For example, for $\boldsymbol{G}\in\mathbb{R}^{I\times J}$, $\boldsymbol{G}[i]$ refers to the $i$th slice along the first dimension, and $\boldsymbol{G}[:,j]$ refers to the $j$th slice along the second dimension, with sizes $\mathbb{R}^{J}$ and $\mathbb{R}^{I}$, respectively. Furthermore, it can also accept a vector $\boldsymbol{v}\in\mathbb{N}^{l}$ as an input. In this case, $\boldsymbol{G}[\boldsymbol{v}]\in\mathbb{R}^{l\times J}$. The Flatten function converts multidimensional input into a one-dimensional vector format. As the LinearSumAssignment function, $\operatorname{scipy.optimize.linear\_sum\_assignment}$ [^1] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.linear_sum_assignment.html is used to solve the LAP. In the AdjustTree function, the parameters of a tree are modified according to the $u$th pattern among the enumerated $U\in\mathbb{N}$ total additional invariances patterns. Additionally, in the Weighting function, parameters are multiplied by the square root of their weights to simulate the process of assessing a rule set. If the first argument for the UpdateBestOperation function, the input inner product, is larger than any previously input inner product values, then $u^{\prime}$ is updated with $u$, the second argument. If not, $u^{\prime}$ remains unchanged.
Complexity. The time complexity of solving the LAP is $\mathcal{O}(M^{3})$ using a modified Jonker-Volgenant algorithm without initialization [25], where $M$ is the number of trees. This process needs to be performed only once in both WM and AM to consider tree permutation invariance. However, the number of additional invariance patterns $U$ scales rapidly as $D$ increases. In a non-oblivious perfect binary tree with depth $D$, there are $12^{D}-$ 1 splitting nodes, resulting in $2^{2^{D}-1}$ possible combinations of sign flips, giving total additional invariances pattern $U=2^{2^{D}-1}$. Additionally, in the case of oblivious trees with depth $D$, the number of splitting rules that consider sign flipping is reduced from $2^{2^{D}-1}$ to $2^{D}$ due to the splitting rule sharing at the same depth. Considering the $D!$ distinct splitting order invariance patterns, we have $U=2^{D}D!$. Therefore, for large values of $D$, it becomes impractical to conduct an exhaustive search to consider additional invariances.
3.3 Architecture-dependency of the Invariances
In Section 3.1, we focused on perfect binary trees as they are most commonly used in soft trees [26, 9, 10]. However, tree architectures can be flexible, and we show that we can specifically design architecture that has neither subtree flip nor splitting order invariances. This allows efficient matching as it is computationally expensive to consider two such invariances.
Table 1: Invariances inherent to each model architecture.
Perm Flip Order
Non-Oblivious Tree $\checkmark$ $\checkmark$ $\times$
Oblivious Tree $\checkmark$ $\checkmark$ $\checkmark$
Decision List $\checkmark$ ()($\checkmark$)( ✓ ) $\times$
Decision List (Modified) $\checkmark$ $\times$ $\times$
4 Experiment
4.1 Setup
Datasets.  In our experiments, we employed Tabular-Benchmark [27], a collection of tabular datasets suitable for evaluating tree ensembles. Details of datasets are provided in Section A in Appendix. As proposed in [27], we randomly sampled $10,000$ instances for train and test data from each dataset. If the dataset contains fewer than $20,000$ instances, they are randomly divided into halves for train and test data. We applied quantile transformation to each feature and standardized it to follow a normal distribution.
Resources.  All experiments were conducted on a system equipped with an Intel Xeon E5-2698 CPU at 2.20 GHz, 252 GB of memory, and Tesla V100-DGXS-32GB GPU, running Ubuntu Linux (version 4.15.0-117-generic). The reproducible PyTorch [29] implementation is provided in the supplementary material.
4.2 Results for Perfect Binary Trees
Figure 6 shows how the barrier between two perfect binary tree model pairs changes in each operation. The vertical axis of each plot in Figure 6 shows the averaged barrier over datasets for each considered invariance. The results for both the oblivious and non-oblivious trees are plotted separately in a vertical layout. The panels on the left display the results when the depth $D$ of the tree varies, keeping $M=256$ constant. The panels on the right show the results when the number of trees $M$ varies, with $D$ fixed at $2$. For both oblivious and non-oblivious trees, we observed that the barrier decreases significantly as the considered invariances increase. Focusing on the test data results, after accounting for various invariances, the barrier is nearly zero, indicating that LMC has been achieved.
Figure 7: Interpolation curves of test accuracy for oblivious trees on 16 datasets from Tabular-Benchmark [27]. Two model pairs are trained on the same dataset. The error bars show the standard deviations of $5$ executions.
In particular, the difference between the case of only permutation and the case where additional invariances are considered tends to be larger in the case of AM. This is because parameter values are not used during the rearrangement of the tree in AM. Additionally, it has been observed that the barrier increases as trees become deeper and that the barrier decreases as the number of trees increases. These behaviors correspond to the changes observed in neural networks when the depth varies or when the width of hidden layers increases [3, 4]. Figure 7 shows interpolation curves for AM in oblivious trees with $D=2$ and $M=256$. In our figures and tables, “Naive” refers to a straightforward parameter interpolation without any specific optimization; “Tree Permutation” or “Perm” considers only the permutation invariance; and “Ours” incorporates both the permutation of the trees and tree-inherent invariances. Other details, such as performance for each dataset, are provided in Section B in Appendix.
Figure 8: Interpolation curves of test accuracy for oblivious trees on 16 datasets from Tabular-Benchmark [27]. Two model pairs are trained on split datasets with different class ratios. The error bars show the standard deviations of $5$ executions. Performance of a model trained with the full dataset is shown in the red dashed horizontal lines as a reference.
Furthermore, we conducted experiments with split data following the protocol in [4] and [30], where the initial split consists of randomly sampled 80% negative and 20% positive instances, and the second split inverts these ratios. There is no overlap between the two split datasets. We trained two model pairs using these separately split datasets and observed an improvement in performance by interpolating their parameters. Figure 8 illustrates the interpolation curves under AM in oblivious trees with parameters $D=2$ and $M=256$. Through model merging, it demonstrates similar performance to full data training even with split data training for the majority of datasets. Note that the data split is configured to remain consistent even when the training random seeds differ. Detailed results for each dataset using WM or AM are provided in Section B in Appendix.
MLP
Barrier
Depth Naive Perm Accuracy Size
1 8.755 ± 0.877 0.491 ± 0.062 76.286 ± 0.094 12034
2 15.341± 1.125 2.997 ± 0.709 75.981 ± 0.139 77826
3 15.915 ± 2.479 5.940 ± 2.153 75.935 ± 0.117 143618
Non-Oblivious Tree
Barrier
Depth Naive Perm Ours Accuracy Size
1 8.965 ± 0.963 0.449 ± 0.235 0.181 ± 0.078 76.464 ± 0.167 12544
2 6.801 ± 0.464 0.811 ± 0.333 0.455 ± 0.105 76.631 ± 0.052 36608
3 5.602 ± 0.926 1.635 ± 0.334 0.740 ± 0.158 76.339 ± 0.115 84736
Oblivious Tree
Barrier
Depth Naive Perm Ours Accuracy Size
1 8.965 ± 0.963 0.449 ± 0.235 0.181 ± 0.078 76.464 ± 0.167 12544
2 7.881 ± 0.866 0.918 ± 0.092 0.348 ± 0.172 76.623 ± 0.042 25088
3 7.096 ± 0.856 1.283 ± 0.139 0.484 ± 0.049 76.535 ± 0.063 38656
Table 2 compares the average test barriers of an MLP with a ReLU activation function, whose width is equal to the number of trees, $M=256$. The procedure for MLPs follows that described in Section 4.1. The permutation for MLPs is optimized using the method described in [4]. Since [4] indicated that WM outperforms AM in neural networks, WM was used for the comparison. Overall, tree models exhibit smaller barriers compared to MLPs while maintaining similar accuracy levels. It is important to note that MLPs with $D>1$ tend to have more parameters at the same depth compared to trees, leading to more complex optimization landscapes. Nevertheless, the barrier for the non-oblivious tree at $D=3$ is smaller than that for the MLP at $D=2$, even with more parameters. Furthermore, at the same depth of $D=1$, tree models have a smaller barrier. Here, the model size is evaluated using $F=44$, the average input feature size of 16 datasets used in the experiments.
In Section 3.2, we have shown that considering additional invariances for deep perfect binary trees is computationally challenging, which may suggest developing heuristic algorithms for deep trees. However, we consider this to be a low priority, supported by our observations that the barrier tends to increase as trees deepen even if we consider invariances. This trend indicates that deep models are fundamentally less important for model merging considerations. Furthermore, deep perfect binary trees are rarely used in practical scenarios. [12] demonstrated that generalization performance degrades with increasing depth in perfect binary trees due to the degeneracy of the Neural Tangent Kernel (NTK) [31]. This evidence further supports the preference for shallow perfect binary trees, and increasing the number of trees can enhance the expressive power while reducing barriers.
4.3 Results for Decision Lists
Table 3: Barriers averaged for 16 datasets under WM with $D=2$ and $M=256$.
Train Test
Barrier Barrier
Architecture Naive Perm Ours Accuracy Naive Perm Ours Accuracy
Table 4: Barriers averaged for 16 datasets under AM with $D=2$ and $M=256$.
Train Test
Barrier Barrier
Architecture Naive Perm Ours Accuracy Naive Perm Ours Accuracy
Figure 9 illustrates barriers as a function of depth, considering only permutation invariance, with $M$ fixed at $256$. In this experiment, we have excluded non-oblivious trees from comparison as the number of their parameters exponentially increases as trees deepen, making them infeasible computation. Our proposed modified decision lists reduce the barrier more effectively than both oblivious trees and the original decision lists. However, the barriers of the modified decision lists are still larger than those obtained by considering additional invariances with perfect binary trees. Tables LABEL:tbl:wm_models and LABEL:tbl:am_models show the averaged barriers for 16 datasets, with $D=2$ and $M=256$. Although the barriers of modified decision lists are small when considering only permutations (Perm), perfect binary trees such as oblivious trees with additional invariances (Ours) exhibit smaller barriers, which supports the validity of using oblivious trees as in [9] and [11]. To summarize, when considering the practical use of model merging, if the goal is to prioritize efficient computation, we recommend using our proposed decision list. Conversely, if the goal is to prioritize barriers, it would be preferable to use perfect binary trees, which have a greater number of invariances that maintain the functional behavior.
5 Conclusion
We have presented the first investigation of LMC for soft tree ensembles. We have identified additional invariances inherent in tree architectures and empirically demonstrated the importance of considering these factors. Achieving LMC is crucial not only for understanding the behavior of non-convex optimization from a learning theory perspective but also for implementing practical techniques such as model merging. By arithmetically combining parameters of differently trained models, a wide range of applications have been explored, such as federated leanning [32] and continual learning [33]. Our research extends these techniques to soft tree ensembles. Future work will explore empirical investigations, including the perspective of general mode connectivity [34].
Acknowledgements
This work was supported by JSPS, KAKENHI Grant Number JP21H03503, Japan and JST, CREST Grant Number JPMJCR22D3, Japan.
Appendix A Dataset
Details of datasets used in experiments are provided in Table 5.